OpenStudy (anonymous):

Evaulate the integral by parts \[\int\limits_{0}^{1}y/e ^{2y} dy\]

OpenStudy (anonymous):

\[\int_0^1 \frac{y}{e^{2y} }dy\] ?

OpenStudy (anonymous):

yah that

OpenStudy (anonymous):

im not getting the correct answer

OpenStudy (anonymous):

the answer is 1/4 - (3/4)(e^-2)

OpenStudy (anonymous):

lets try this again \[\int ye^{-2y}dy\] \[\frac{y e^{-2y}}{-2} - \int\frac{e^{-2y}}{-2}\] \[ - \frac{ye^{-2y}}{2} - \frac{e^{-2y}}{4}\]

OpenStudy (anonymous):

so now when we put the limits

OpenStudy (anonymous):

i did things wrong in my first attempt not this one atleast thats what i think lol \[-\frac{e^{-2}}{2} - \frac{e^{-2}}{4} + \frac{1}{4}\]

OpenStudy (anonymous):

\[\frac{1}{4} - \frac{-3e^{-2}}{4}\]

OpenStudy (anonymous):

ok i kinda get the steps now

OpenStudy (anonymous):

good : )

OpenStudy (anonymous):

gotta change to e^-2y

OpenStudy (anonymous):

yeah that makes it easy to think and perform in my first attempt i didn't do it and had trouble

OpenStudy (anonymous):

yeah thats what i didnt do